# Learning Time-Varying Multi-Region Communications via Scalable Markovian Gaussian Processes

### v2

@@ 4.1 Synthetic Data
In this section, we simulate a common phenomenon in neuroscience where brain region communications are dynamic Parra & Tobar (2017). Our goal is to evaluate our model’s ability to recover time-varying temporal delays and latent variables.
@@ 4.2 Two Brain Regions
We use smoothed multi-region spike trains from session 106r001p26 with an orientation of $0^{\circ}$. This dataset consists of 400 trials, each containing 64 time bins (20 ms per bin), with 72 V1 neurons and 22 V2 neurons. The monkey begins receiving the visual stimulus (drifting gratings) at the first time bin, and the stimulus persists throughout all 64 time bins. The number of across-region and within-region latent dynamics follows previous works Gokcen et al. (2022); Li et al. (2024), where $m_{a}=2$ and $m_{w}=2$. The order $P=4$ is selected based on performance evaluation on the validation dataset.
@@ Results.
Figure 2(A) presents the estimated across-region communications and time-varying temporal delays from the test dataset (showing ten trials, see Appendix E for within-region dynamics). Both groups of across-region communications exhibit a periodic pattern driven by the external drifting grating stimulus. Our model reveals a slow forward and feedback communication circuit (characterized by larger delays) between V1 and V2 immediately after stimulus onset. Over time, this shifts to a dominance of faster forward communication (with smaller delays). The time-varying temporal delays can be interpreted as follows: shortly after stimulus presentation, V1 triggers oscillatory dynamics in V2, which in turn generates a strong feedback signal from V2 to V1, potentially reflecting the emergence of surprise or prediction error Rao & Ballard (1999). Subsequently, both regions begin to oscillate synchronously with dominant forward communication. Our findings on V1-V2 interactions align with previous studies Gokcen et al. (2022); Li et al. (2024); Gokcen et al. (2024a).
Figure 2(B) compares the test log-likelihood summed over trials and time bins, showing that our model, ADM, outperforms MRM-GP and DLAG, attributed to its ability to capture continuously time-varying temporal delays. Figure 2(C) presents the validation log-likelihood across different $P$ values, with $P=4$ achieving the highest value. Combined with the insights from Figures 2(A–B), this suggests that a small $P$ value can effectively estimate model parameters and latent variables.
@@ 4.3 Five Brain Regions
Figure 3: Evaluation of the ADM on spike trains from five visual brain regions. (A) Estimated pairwise temporal delays from one group of across-region communications, showing time-varying dynamics and alignment with the anatomical hierarchy of the mouse visual cortex. (B) Meso-scale brain networks at $t=3$ and $t=50$ time bins, illustrating changes in communication speed and direction. (C) Cross-validation results for hyperparameter selection, with $m_{a}=3$, $m_{w}=1$, and $P=2$ achieving the highest validation log-likelihood. (D) Test log-likelihood comparison between ADM and mDLAG, demonstrating ADM’s superior fit due to its ability to model time-varying communications.
In this section, we scale up our model to a larger neural recording spanning five regions with increased time resolution. Our objective is to investigate across-region communications and identify the time-varying meso-scale brain network, defined as the dynamic network spanning sub-brain regions, e.g., regions in visual cortex.
@@
We use smoothed multi-region spike train data from the Visual Coding – Neuropixels project by the Allen Institute Siegle et al. (2021), specifically from session 750749662. This dataset includes spike trains recorded from VISp, VISrl, VISal, VISpm, and VISam—sub-areas of the mouse visual cortex. It consists of 120 trials, $T=200$ time bins (each 10 ms), and a total of 202 neurons, with external visual stimuli comprising 4 Hz drifting gratings. Following the approach in Gokcen et al. (2022), we first apply Factor Analysis to estimate the total number of across-region and within-region latent dynamics, determining $M=4$ (see Appendix F for details). We then conduct a grid search with 5-fold cross-validation to refine the number of across-region and within-region latent dynamics and the model order $P$, yielding $m_{a}=3$, $m_{w}=1$, and $P=2$.
@@ Results.
Figure 3(A) presents the ten estimated pairwise temporal delays from one group of across-region communications. For the remaining groups and estimated latent dynamics, see Appendix E. Our results reveal consistent forward communication from VISp to downstream visual areas, such as VISrl, VISal, and VISpm, aligning with the known anatomical hierarchy of the mouse visual cortex Siegle et al. (2021). Additionally, these forward communications exhibit time-varying dynamics. For instance, communication between VISp and VISrl transitions from slow to fast, indicating an enhanced interaction that gradually becomes more synchronous following the initial surprise response to the visual stimulus onset. In contrast, the communication between VISp and VISal shifts from fast to slow, suggesting inhibition induced by the external stimulus. Furthermore, our findings indicate that all communications involving VISam are feedback signals. This is expected, as VISam is positioned at the end of the anatomical hierarchy of the mouse visual system, consistent with the anatomical hierarchy scores reported in Siegle et al. (2021).
Figure 3(B) illustrates the corresponding meso-scale brain network based on the estimated temporal delays. Each node represents a region in the visual system, while directed edges indicate the directional communications. The length of each edge is determined by the estimated delays, reflecting the speed of communication. The figure presents two meso-scale brain networks at $t=3$ and $t=50$ time bins. The primary differences between these networks include changes in the speed of certain forward communications and a shift in the communication direction between VISrl and VISal, suggesting the emergence of gradually synchronous interactions following stimulus presentation.
Figure 3(C) presents the cross-validation results for twenty hyperparameter combinations. We first determine $M=4$ using Factor Analysis (see Appendix F for details), then conduct a grid search over all combinations of $m_{a}\in[0,4]$, $m_{w}\in[0,4]$, and $P\in[2,5]$. The highest validation log-likelihood is achieved with $m_{a}=3$, $m_{w}=1$, and $P=2$.
Figure 3(D) compares the test log-likelihood, summed over trials and time bins, between our model (ADM) and mDLAG, which is an extension of DLAG that supports more than two brain regions using a variational inference method. The results indicate that ADM provides a better fit to the data, attributed to its ability to model time-varying communications. We do not compare MRM-GP and DLAG since they are limited to two brain regions. Additionally, we skip a time cost comparison with mDLAG because it is implemented only in MATLAB, which is significantly slower (taking hours to run) than our GPU-optimized implementation.
@@ 5 Discussion
Let’s inspect $\mathbf{V}\mathbf{V}^{\top}\in\mathbb{R}^{NP\times NP}$ first. We have:
$\begin{split}\mathbf{V}\mathbf{V}^{\top}&=\begin{bmatrix}x_{1}&x_{2}&\dots&x_{T-P}\\ x_{2}&x_{3}&\dots&x_{T-P+1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{P}&x_{P+1}&\dots&x_{T-1}\end{bmatrix}\begin{bmatrix}x_{1}^{\top}&x_{2}^{\top}&\dots&x_{P}^{\top}\\ x_{2}^{\top}&x_{3}^{\top}&\dots&x_{P+1}^{\top}\\ \vdots&\vdots&\ddots&\vdots\\ x_{T-P}^{\top}&x_{T-P+1}^{\top}&\dots&x_{T-1}^{\top}\end{bmatrix}\\ &=\begin{bmatrix}x_{1}x_{1}^{\top}+\dots+x_{T-P}x_{T-P}^{\top}&\dots&x_{1}x_{P}^{\top}+x_{2}x_{P+1}^{\top}+\dots+x_{T-P}x_{T-1}^{\top}\\ \vdots&\ddots&\vdots\\ x_{P}x_{1}^{\top}+\dots+x_{T-1}x_{T-P}^{\top}&\dots&x_{P}x_{P}^{\top}+x_{P+1}x_{P+1}^{\top}+\dots+x_{T-1}x_{T-1}^{\top}\\ \end{bmatrix},\end{split}$ (15)
where the first element $x_{1}x_{1}^{\top}\in\mathbb{R}^{N\times N}$ represents the auto-covariance of $x_{1}$, which is essentially the kernel $\mathbf{K}(0)$ (Eq. 4). In other words, since $\boldsymbol{x}$ is modeled as a stationary GP, the elements $x_{1}x_{1}^{\top},\dots,x_{T-1}x_{T-1}^{\top}$ are all equivalent and correspond to the diagonal elements $\mathbf{K}(0)$ in Eq. 4. Similarly, the elements $x_{P}x_{1}^{\top},\dots,x_{T-1}x_{T-P}^{\top}$ represent cross-covariances with time interval $P-1$, which correspond to the off-diagonal elements $\mathbf{K}(P-1)$. Therefore, we can further write Eq. 15 as:
@@ Appendix B Details for Kalman EM
In E-step, we have the expected log-likelihood as:
$Q(\Theta|\Theta^{k})$ $=$ $\mathbb{E}_{\boldsymbol{x}\sim p(\cdot|\boldsymbol{y},\Theta^{k})}[LL(\boldsymbol{x},\boldsymbol{y}|\Theta)],$
@@
$\propto$ $-\frac{1}{2}\sum_{t}(y_{t}^{\top}\mathbf{C}^{-1}y_{t}-2y_{t}^{\top}\mathbf{C}^{-1}\mathbf{H}\mathbb{E}[x_{t}|\boldsymbol{y},\Theta^{k}]+\mathbb{E}[x_{t}^{\top}\mathbf{H}^{\top}\mathbf{C}^{-1}\mathbf{H}x_{t}|\boldsymbol{y},\Theta^{k}])$
$-\frac{1}{2}\sum_{t}(\mathbb{E}[x_{t}^{\top}\mathbf{\hat{Q}}^{-1}x_{t}|\boldsymbol{y},\Theta^{k}]-2\mathbb{E}[x_{t}^{\top}\mathbf{\hat{Q}}^{-1}\mathbf{\hat{A}}x_{t-1}|\boldsymbol{y},\Theta^{k}]$
$+\mathbb{E}[x_{t-1}^{\top}\mathbf{\hat{A}}^{\top}\mathbf{\hat{Q}}^{-1}\mathbf{\hat{A}}x_{t-1}|\boldsymbol{y},\Theta^{k}])-\frac{T}{2}\log(|\mathbf{C}|)-\frac{1}{2}\log(|\mathbf{\hat{Q}}|),$ (20)
where all expectations $\mathbb{E}[\cdot]$ mentioned above are based on $\mathbb{E}[x_{t}|\boldsymbol{y}]$, $\mathbb{E}[x_{t}^{\top}x_{t}|\boldsymbol{y}]$, and $\mathbb{E}[x_{t}^{\top}x_{t-1}|\boldsymbol{y}]$, which are calculated through forward and backward passes known as Kalman Filter and Kalman Smoother Boots (2009), respectively. Specifically, if denoting $\widetilde{x_{t}}=\mathbb{E}[x_{t}|\boldsymbol{y}]$, $\mathbf{C}_{t}=\mathbb{E}[x_{t}^{\top}x_{t}|\boldsymbol{y}]$, and $\mathbf{C}_{t,t-1}=\mathbb{E}[x_{t}^{\top}x_{t-1}|\boldsymbol{y}]$, we have:
$\begin{split}\mathbb{E}&[x_{t}|\boldsymbol{y},\Theta^{k}]=\widetilde{x_{t}},\\ \mathbb{E}&[x_{t}^{\top}\mathbf{H}^{\top}\mathbf{C}^{-1}\mathbf{H}x_{t}|\boldsymbol{y},\Theta^{k}]=\widetilde{x_{t}}^{\top}\mathbf{H}^{\top}\mathbf{C}^{-1}\mathbf{H}\widetilde{x_{t}}+\operatorname{Tr}(\mathbf{H}^{\top}\mathbf{C}^{-1}\mathbf{H}\mathbf{C}_{t}),\\ \mathbb{E}&[x_{t}^{\top}\mathbf{\hat{Q}}^{-1}x_{t}|\boldsymbol{y},\Theta^{k}]=\widetilde{x_{t}}^{\top}\mathbf{\hat{Q}}^{-1}\widetilde{x_{t}}+\operatorname{Tr}(\mathbf{\hat{Q}}^{-1}\mathbf{C}_{t}),\\ \mathbb{E}&[x_{t}^{\top}\mathbf{\hat{Q}}^{-1}\mathbf{\hat{A}}x_{t-1}|\boldsymbol{y},\Theta^{k}]=\widetilde{x_{t}}^{\top}\mathbf{\hat{Q}}^{-1}\mathbf{\hat{A}}\widetilde{x}_{t-1}+\operatorname{Tr}(\mathbf{\hat{Q}}^{-1}\mathbf{\hat{A}}\mathbf{C}_{t,t-1}^{\top}),\\ \mathbb{E}&[x_{t-1}^{\top}\mathbf{\hat{A}}^{\top}\mathbf{\hat{Q}}^{-1}\mathbf{\hat{A}}x_{t-1}|\boldsymbol{y},\Theta^{k}]=\widetilde{x}_{t-1}^{\top}\mathbf{\hat{A}}^{\top}\mathbf{\hat{Q}}^{-1}\mathbf{\hat{A}}\widetilde{x}_{t-1}+\operatorname{Tr}(\mathbf{\hat{A}}^{\top}\mathbf{\hat{Q}}^{-1}\mathbf{\hat{A}}\mathbf{C}_{t}).\end{split}$ (21)
Then, in M-step, we aim to update the kernel parameters such that $Q(\Theta|\Theta^{k})$ is maximized:
[…]

### v4

@@ 4.1 Synthetic Data
In this section, we simulate a common phenomenon in neuroscience where brain region communications are dynamic (Parra & Tobar, 2017). Our goal is to evaluate our model’s ability to recover time-varying temporal delays and latent variables.
@@ Results.
Finally, Appendix G presents a more complex case with five-region synthetic dataset and evaluates the model’s performance as a function of the number of regions, latent variables, and data length.
@@ 4.2 Two Brain Regions
We use smoothed multi-region spike trains from session 106r001p26 with an orientation of $0^{\circ}$. This dataset consists of 400 trials, each containing 64 time bins (20 ms per bin), with 72 V1 neurons and 22 V2 neurons. The monkey begins receiving the visual stimulus (drifting gratings) at the first time bin, and the stimulus persists throughout all 64 time bins. The number of across-region and within-region latent dynamics follows previous works (Gokcen et al., 2022; Li et al., 2024), where $m_{a}=2$ and $m_{w}=2$. The order $P=4$ is selected based on performance evaluation on the validation dataset.
@@ Results.
Figure 2(A) shows the estimated across-region communications and time-varying delays from the test dataset (ten trials shown; within-region dynamics are provided in Appendix E). The first group of communications shows a shift from slower feedback (with larger absolute delays) to faster feedback (with smaller absolute delays) starting around 200 ms after stimulus onset. In contrast, the second group exhibits a periodic pattern driven by the external drifting grating stimulus, along with a change in communication direction immediately following stimulus onset. The time-varying delays suggest the following interpretation: shortly after stimulus presentation, V2 generates a strong feedback signal from V2 to V1, potentially reflecting the emergence of surprise or prediction error (Rao & Ballard, 1999). As time progresses, both regions transition into more synchronized oscillations. Our findings on V1-V2 interactions are similar to previous studies (Gokcen et al., 2022; Li et al., 2024; Gokcen et al., 2024a).
Figure 2(B) compares the test log-likelihood summed over trials and time bins, showing that our model, ADM, outperforms MRM-GP and DLAG under the same $m_{a}$ and $m_{w}$ settings. This improvement is attributed to ADM’s ability to capture continuously time-varying temporal delays. Figure 2(C) presents the validation log-likelihood across different $P$ values, with $P=4$ achieving the highest value. Combined with the insights from Figures 2(A–B), this suggests that a small $P$ value can effectively estimate model parameters and latent variables.
@@ 4.3 Five Brain Regions
Figure 3: Evaluation of the ADM on spike trains from five visual brain regions. (A) Estimated pairwise temporal delays from one group of across-region communications, showing time-varying dynamics and alignment with the anatomical hierarchy of the mouse visual cortex. (B) Meso-scale brain networks at $t=3$ and $t=50$ time bins, derived from the communications in (A), illustrating changes in communication speed and direction. (C) Cross-validation results for hyperparameter selection, with $m_{a}=3$, $m_{w}=1$, and $P=5$ achieving the highest validation log-likelihood. (D) Test log-likelihood comparison between ADM and mDLAG, demonstrating ADM’s superior fit due to its ability to model time-varying communications.
In this section, we scale up our model to a larger neural recording spanning five regions with increased time resolution. Our objective is to investigate across-region communications and identify the time-varying meso-scale brain network, defined as the dynamic network spanning sAppendixub-brain regions, e.g., regions in visual cortex.
@@
We use smoothed multi-region spike train data from the Visual Coding – Neuropixels project by the Allen Institute (Siegle et al., 2021), specifically from session 750749662. This dataset includes spike trains recorded from VISp, VISrl, VISal, VISpm, and VISam—sub-areas of the mouse visual cortex. It consists of 120 trials, $T=200$ time bins (each 10 ms), and a total of 202 neurons, with external visual stimuli comprising 4 Hz drifting gratings. Following the approach in (Gokcen et al., 2022), we first apply Factor Analysis to estimate the total number of across-region and within-region latent dynamics, determining $M=4$ (see Appendix F for details). We then conduct a grid search with 5-fold cross-validation to refine the number of across-region and within-region latent dynamics and the model order $P$.
@@ Results.
Figure 3(A) presents the ten estimated pairwise temporal delays from one group of across-region communications. See Appendix E for the estimated latent dynamics. Our results reveal consistent forward communication from VISp to downstream visual areas, such as VISrl, VISal, and VISpm, aligning with the known anatomical hierarchy of the mouse visual cortex (Siegle et al., 2021). Additionally, these forward communications exhibit time-varying dynamics. For instance, communication between VISp and VISrl transitions from slow to fast, indicating an enhanced interaction that gradually becomes more synchronous following the initial surprise response to the visual stimulus onset. In contrast, the communication between VISp and VISal shifts from fast to slow, suggesting inhibition induced by the external stimulus. Furthermore, our findings indicate that all communications involving VISam are feedback signals. This is expected, as VISam is positioned at the end of the anatomical hierarchy of the mouse visual system, consistent with the anatomical hierarchy scores reported in (Siegle et al., 2021).
Figure 3(B) depicts the meso-scale brain network corresponding to the across-region communications presented in Figure 3(A). Each node represents a region in the visual system, while directed edges indicate the directional communications. The length of each edge is determined by the estimated delays, reflecting the speed of communication. The figure presents two meso-scale brain networks at $t=3$ and $t=50$ time bins. The primary differences between these networks include changes in the speed of certain forward communications and a direction change in the communication direction between VISrl and VISal, suggesting the emergence of stimulus presentation.
Figure 3(C) presents the cross-validation results for twenty hyperparameter combinations. We first determine $M=4$ using Factor Analysis (see Appendix F for details), then conduct a grid search over all combinations of $m_{a}\in[0,4]$, $m_{w}\in[0,4]$, and $P\in[2,5]$. The highest validation log-likelihood is achieved with $m_{a}=3$, $m_{w}=1$, and $P=5$.
Figure 3(D) compares the test log-likelihood, summed over trials and time bins, between our model (ADM) and mDLAG with $m_{a}=3$ latent communication channels, where mDLAG is an extension of DLAG that supports more than two brain regions using variational inference. The results indicate that ADM provides a better fit to the data, attributed to its ability to model time-varying communications. We do not compare MRM-GP and DLAG since they are limited to two brain regions. Additionally, we skip a time cost comparison with mDLAG because it is implemented only in MATLAB, which is significantly slower than our GPU-optimized implementation.
@@ 5 Discussion
Acknowledgement
This work is supported by National Institutes of Health BRAIN initiative (1U01NS131810).
Impact Statement
Our work presents the Adaptive Delay Model (ADM), a scalable and biologically relevant framework for uncovering time-varying multi-region neural communications. By bridging Gaussian Processes (GPs) with State Space Models (SSMs), ADM enhances our ability to analyze large-scale neural recordings, offering insights into sensory processing, cognitive flexibility, and neural disorders. Beyond neuroscience, its adaptability to fields such as robotics, autonomous systems, and signal processing broadens its societal impact. Ethically, as machine learning models increasingly shape neuroscience research, it is crucial to ensure their responsible application, avoiding overinterpretation of inferred neural dynamics and considering the broader implications for brain-computer interfaces and neurotechnology.
@@ Acknowledgement
Let’s inspect $\mathbf{V}\mathbf{V}^{\top}\in\mathbb{R}^{NP\times NP}$ first. To simplify the notation, we use $\boldsymbol{x}$ to represent $\boldsymbol{x}_{m}^{a}$ in Eq. 7. We have:
$\begin{split}\mathbf{V}\mathbf{V}^{\top}&=\begin{bmatrix}\boldsymbol{x}_{1}&\boldsymbol{x}_{2}&\dots&\boldsymbol{x}_{T-P}\\ \boldsymbol{x}_{2}&\boldsymbol{x}_{3}&\dots&\boldsymbol{x}_{T-P+1}\\ \vdots&\vdots&\ddots&\vdots\\ \boldsymbol{x}_{P}&\boldsymbol{x}_{P+1}&\dots&\boldsymbol{x}_{T-1}\end{bmatrix}\begin{bmatrix}\boldsymbol{x}_{1}^{\top}&\boldsymbol{x}_{2}^{\top}&\dots&\boldsymbol{x}_{P}^{\top}\\ \boldsymbol{x}_{2}^{\top}&\boldsymbol{x}_{3}^{\top}&\dots&\boldsymbol{x}_{P+1}^{\top}\\ \vdots&\vdots&\ddots&\vdots\\ \boldsymbol{x}_{T-P}^{\top}&\boldsymbol{x}_{T-P+1}^{\top}&\dots&\boldsymbol{x}_{T-1}^{\top}\end{bmatrix}\\ &=\begin{bmatrix}\boldsymbol{x}_{1}\boldsymbol{x}_{1}^{\top}+\dots+\boldsymbol{x}_{T-P}\boldsymbol{x}_{T-P}^{\top}&\dots&\boldsymbol{x}_{1}\boldsymbol{x}_{P}^{\top}+\boldsymbol{x}_{2}\boldsymbol{x}_{P+1}^{\top}+\dots+\boldsymbol{x}_{T-P}\boldsymbol{x}_{T-1}^{\top}\\ \vdots&\ddots&\vdots\\ \boldsymbol{x}_{P}\boldsymbol{x}_{1}^{\top}+\dots+\boldsymbol{x}_{T-1}\boldsymbol{x}_{T-P}^{\top}&\dots&\boldsymbol{x}_{P}\boldsymbol{x}_{P}^{\top}+\boldsymbol{x}_{P+1}\boldsymbol{x}_{P+1}^{\top}+\dots+\boldsymbol{x}_{T-1}\boldsymbol{x}_{T-1}^{\top}\\ \end{bmatrix},\end{split}$ (15)
where the first element $\boldsymbol{x}_{1}\boldsymbol{x}_{1}^{\top}\in\mathbb{R}^{N\times N}$ represents the auto-covariance of $\boldsymbol{x}_{1}$, which is essentially the kernel $\mathbf{K}(0)$ (Eq. 4). In other words, since $\boldsymbol{x}$ is modeled as a stationary GP, the elements $\boldsymbol{x}_{1}\boldsymbol{x}_{1}^{\top},\dots,\boldsymbol{x}_{T-1}\boldsymbol{x}_{T-1}^{\top}$ are all equivalent and correspond to the diagonal elements $\mathbf{K}(0)$ in Eq. 4. Similarly, the elements $\boldsymbol{x}_{P}\boldsymbol{x}_{1}^{\top},\dots,\boldsymbol{x}_{T-1}\boldsymbol{x}_{T-P}^{\top}$ represent cross-covariances with time interval $P-1$, which correspond to the off-diagonal elements $\mathbf{K}(P-1)$. Therefore, we can further write Eq. 15 as:
@@ Appendix B Details for Kalman EM
We consider the linear–Gaussian state–space model
$\boldsymbol{x}_{t}=\mathbf{A}_{t}\boldsymbol{x}_{t-1}+\boldsymbol{w}_{t},\qquad\boldsymbol{w}_{t}\sim\mathcal{N}\!\bigl{(}0,\mathbf{Q}_{t}\bigr{)},\qquad\boldsymbol{y}_{t}=\mathbf{H}\boldsymbol{x}_{t}+\boldsymbol{v}_{t},\qquad\boldsymbol{v}_{t}\sim\mathcal{N}\!\bigl{(}0,\mathbf{V}\bigr{)},$
where the transition matrix $\mathbf{A}_{t}$ and process-noise covariance $\mathbf{Q}_{t}$ change with time. Let $\Theta=\{\mathbf{A}_{1:T},\mathbf{Q}_{1:T},\mathbf{H},\mathbf{V}\}$ denote the full parameter set.
E–step
With parameters fixed at $\Theta^{k}$, the expected complete-data log-likelihood is
$Q(\Theta\mid\Theta^{k})$ $=\mathbb{E}_{\mathbf{x}\mid\mathbf{y},\Theta^{k}}\!\bigl{[}\log p(\mathbf{x},\mathbf{y}\mid\Theta)\bigr{]}$
@@
$\propto-\frac{1}{2}\sum_{t=1}^{T}\!\Bigl{(}\boldsymbol{y}_{t}^{\!\top}\mathbf{V}^{-1}\boldsymbol{y}_{t}-2\boldsymbol{y}_{t}^{\!\top}\mathbf{V}^{-1}\mathbf{H}\,\widetilde{\mathbf{x}}_{t}+\operatorname{Tr}\!\bigl{(}\mathbf{H}^{\top}\mathbf{V}^{-1}\mathbf{H}\mathbf{V}_{t}\bigr{)}\Bigr{)}$
$\hphantom{\propto}\;-\frac{1}{2}\sum_{t=1}^{T}\!\Bigl{(}\operatorname{Tr}\!\bigl{(}\mathbf{Q}_{t}^{-1}\mathbf{C}_{t}\bigr{)}-2\operatorname{Tr}\!\bigl{(}\mathbf{Q}_{t}^{-1}\mathbf{A}_{t}\mathbf{C}_{t,t-1}^{\!\top}\bigr{)}+\operatorname{Tr}\!\bigl{(}\mathbf{A}_{t}^{\!\top}\mathbf{Q}_{t}^{-1}\mathbf{A}_{t}\mathbf{C}_{t-1}\bigr{)}\Bigr{)}$
$\hphantom{\propto}\;-\frac{T}{2}\log|\mathbf{V}|-\frac{1}{2}\sum_{t=1}^{T}\log|\mathbf{Q}_{t}|.$ (20)
Here
$\widetilde{\mathbf{x}}_{t}=\mathbb{E}[\boldsymbol{x}_{t}\mid\mathbf{y},\Theta^{k}],\quad\mathbf{C}_{t}=\mathbb{E}[\boldsymbol{x}_{t}\boldsymbol{x}_{t}^{\!\top}\mid\mathbf{y},\Theta^{k}],\quad\mathbf{C}_{t,t-1}=\mathbb{E}[\boldsymbol{x}_{t}\boldsymbol{x}_{t-1}^{\!\top}\mid\mathbf{y},\Theta^{k}],$
which are obtained with a Kalman filter followed by a Rauch–Tung–Striebel smoother run with the time-varying parameters (Boots, 2009). The individual expectations that appear in (20) expand to
$\mathbb{E}[\boldsymbol{x}_{t}^{\!\top}\mathbf{H}^{\top}\mathbf{V}^{-1}\mathbf{H}\boldsymbol{x}_{t}]$ $=\widetilde{\mathbf{x}}_{t}^{\!\top}\mathbf{H}^{\top}\mathbf{V}^{-1}\mathbf{H}\widetilde{\mathbf{x}}_{t}+\operatorname{Tr}\!\bigl{(}\mathbf{H}^{\top}\mathbf{V}^{-1}\mathbf{H}\mathbf{C}_{t}\bigr{)},$ (21)
$\mathbb{E}[x_{t}^{\!\top}\mathbf{Q}_{t}^{-1}\boldsymbol{x}_{t}]$ $=\widetilde{\mathbf{x}}_{t}^{\!\top}\mathbf{Q}_{t}^{-1}\widetilde{\mathbf{x}}_{t}+\operatorname{Tr}\!\bigl{(}\mathbf{Q}_{t}^{-1}\mathbf{C}_{t}\bigr{)},$ (22)
$\mathbb{E}[\boldsymbol{x}_{t}^{\!\top}\mathbf{Q}_{t}^{-1}\mathbf{A}_{t}x_{t-1}]$ $=\widetilde{\mathbf{x}}_{t}^{\!\top}\mathbf{Q}_{t}^{-1}\mathbf{A}_{t}\widetilde{\mathbf{x}}_{t-1}+\operatorname{Tr}\!\bigl{(}\mathbf{Q}_{t}^{-1}\mathbf{A}_{t}\mathbf{C}_{t,t-1}^{\!\top}\bigr{)},$ (23)
$\mathbb{E}[\boldsymbol{x}_{t-1}^{\!\top}\mathbf{A}_{t}^{\!\top}\mathbf{Q}_{t}^{-1}\mathbf{A}_{t}x_{t-1}]$ $=\widetilde{\mathbf{x}}_{t-1}^{\!\top}\mathbf{A}_{t}^{\!\top}\mathbf{Q}_{t}^{-1}\mathbf{A}_{t}\widetilde{\mathbf{x}}_{t-1}+\operatorname{Tr}\!\bigl{(}\mathbf{A}_{t}^{\!\top}\mathbf{Q}_{t}^{-1}\mathbf{A}_{t}\mathbf{C}_{t-1}\bigr{)}.$ (24)
M–step
The M–step maximises (20) with respect to $\Theta$,
$\Theta^{k+1}=\operatorname*{arg\,max}_{\Theta}Q(\Theta\mid\Theta^{k}),$
where the paramerer sets $\Theta$ is updated either via gradient descent or in closed form, depending on the specific component.